4. What mass of CuI is present in a 1.2 L saturated solution at 25°C? Ksp = 1.3 x10-12 (5 marks)



5. Will a precipitate form when 125 mL of 4.0 x 10-2 M of CaCl2 is added to 175 mL of 2.9 x 10-2 M of NaOH? Ksp of Ca(OH)2 = 4.8 x 10-6. Show the dissociation equation for the possible precipitate and all formulas. (10 marks)


1
Expert's answer
2022-05-06T15:41:03-0400

5.Ca(OH)2 <---> Ca2+ + 2OH-

Ksp = 4.8•10-6 = [Ca2+][OH-]2

And thus we must work out the concentrations with respect to Ca2+and OH- and determine the initial ion product........Q

[Ca2+] = Moles of calcium ion / volume of solution = (0.125 L • 4.0 x 10-2 M) / (0.125 L + 0.175 L) = 0.117 M


[OH-]2 = Moles of hydroxide ion / volume of solution = (0.175 L • 2.9 x 10-2 M)/(0.125 L + 0.175 L) = 0.017 M

And the ion product 

Q = [Ca2+][OH-]2 = 0.117 • (0.017)2 = 3.38x10-5

Since Ksp > Q, precipitation SHOULD NOT occur